OpenStudy (anonymous):

HELP GIVING MEDALS Find the area of the triangle with the given measurements. Round the solution to the nearest hundredth if necessary. B = 104°, a = 11 cm, c = 18 cm 192.12 cm2 96.06 cm2 23.95 cm2 99 cm2

OpenStudy (anonymous):

http://openstudy.com/study#/updates/518ebc41e4b0cf6dd8f4502c Look at this for how to do it. Could you please fan me?

OpenStudy (danjs):

ILL FAN YOU

OpenStudy (anonymous):

Really? Thanks!

OpenStudy (anonymous):

I fanned you

OpenStudy (danjs):

that was the best answer ever, no work required, simply look

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

@DanJS no bc that answer comes out to 7 something

OpenStudy (danjs):

oh , i didnt read it

OpenStudy (danjs):

um, one sec

OpenStudy (danjs):



OpenStudy (danjs):

find the height h

OpenStudy (danjs):

sin(76) = h/11

OpenStudy (danjs):

then just do area of the triangle with base 18 and height 11*sin(76)

OpenStudy (danjs):

what ya figure..

OpenStudy (anonymous):

sin 76= 0.56610763689

OpenStudy (anonymous):

from your solution ans is 56.044656052 but this is not in option

OpenStudy (danjs):

\[Area = \frac{ 1 }{ 2 }*base*height = \frac{ 1 }{ 2 }*18*11*\sin(76) \approx\]

OpenStudy (danjs):

the answer is there, which part you having trouble calculating

OpenStudy (danjs):

you in degree mode?

OpenStudy (anonymous):

ya i do the same but the ans is wrong

OpenStudy (danjs):

96.0593

OpenStudy (anonymous):

plz try your self

OpenStudy (anonymous):

i got like 60

OpenStudy (danjs):

\[0.5*18*11*(0.970296)\]

OpenStudy (danjs):

not sure what ya did

OpenStudy (danjs):

If you got 56. something, you are in radian mode not degree mode

OpenStudy (anonymous):

oh ok thanks so much